OpenStudy (anonymous):

Trini's car breaks down on a highway. Trini estimates that she is 20 to 30 miles from the nearest car repair shop. She calls a towing company that charges a fee of $80 plus $3 per mile to tow a car. If Trini uses this towing company, which is the best estimate for the amount of money, m, she will pay for the company to tow her car? 103 ≤ m ≤ 113 140 ≤ m ≤ 150 140 ≤ m ≤ 170 160 ≤ m ≤ 170

OpenStudy (anonymous):

The amount Trini would have to pay depends on the distance: $80 up front, then $3 times the number of miles traveled, or m=80+3d. The smallest value for d is 20, so the smallest value for m = 80 + 3(20)=140 The largest value for d is 30, so the largest value for m = 80 + 3(30) = 170. So, \[140\le m \le 170\]

OpenStudy (anonymous):

thank you!